# AutoCV: Empowering Reasoning with Automated Process Labeling via Confidence Variation

### v1

@@ Abstract
In this work, we propose a novel method named Automated Process Labeling via Confidence Variation (AutoCV) to enhance the reasoning capabilities of large language models (LLMs) by automatically annotating the reasoning steps. Our approach begins by training a verification model on the correctness of final answers, enabling it to generate automatic process annotations. This verification model assigns a confidence score to each reasoning step, indicating the probability of arriving at the correct final answer from that point onward. We detect relative changes in the verification’s confidence scores across reasoning steps to automatically annotate the reasoning process. This alleviates the need for numerous manual annotations or the high computational costs associated with model-induced annotation approaches. We experimentally validate that the confidence variations learned by the verification model trained on the final answer correctness can effectively identify errors in the reasoning steps. Subsequently, we demonstrate that the process annotations generated by AutoCV can improve the accuracy of the verification model in selecting the correct answer from multiple outputs generated by LLMs. Notably, we achieve substantial improvements across five datasets in mathematics and commonsense reasoning. The source code of AutoCV is available at https://github.com/rookie-joe/AUTOCV.
@@ 3 Method
In the following, we first introduce the problem setting in Sec. Section 3.1. Then we discuss the motivation behind why we believe it is necessary to train a verification model in Section 3.2. Finally, we describe how we transition from outcome supervision to process supervision during the training of the verification model in Section 3.3.
@@ 3.2 Motivation
Table 1: Performance of Mixtral-Instruct on the GSM8K test set. All results are reported in accuracy (%).
@@
Table 2: Comparison of different selection methods across various model sizes (from 7B to 72B) for selecting a response from candidate responses generated by Mixtral-Instruct. All results are reported in Accuracy (%).
@@ 4.1 Experiment on Outcome Supervised Verifier Performance
To evaluate the OSV, we measure its ability to select a sample with the correct final answer from samples generated by various LLMs, denoted as the Response Generator. Specifically, we assess the effectiveness of outcome supervision on two models: Phi2 (OSV-Phi[36] and Mistral-7B (OSV- Mistral[33]. To explore the scalability of this outcome-supervised verifier effect, we choose Response Generators of varying scales, ranging from 7B to 72B parameters, i.e., Mistral-7B-Instruct (Mistral-7B[33], Mixtral 8 $\times$ 7B (Mixtral-8$\times$7B[32] and Qwen-72B-Chat (Qwen-72B[35]. We employ the metric pass@$k$ [37] to assess the performance of those response generators, defined as the scenario where at least one correct instance is within the model’s first $k$ attempts. We then evaluate the OSV’s generalized ranking capability across different LLM scales.
Table 3: Performance of OSV models across different configurations.
@@
The performance disparity among the verifiers can be attributed primarily to variations in model sizes and the accuracy of their training data. Table 4 presents a consolidated view of the model sizes along with the accuracy metrics of their outcome supervision training data.
@@
In our content, we select the OSV (Mistral) model as the OSV model among other experiment settings due to its superior performance, as demonstrated in Table 3.
@@ 5.1 Experimental Setup
Evaluation: We follow the evaluation method in [42], employing the “Accuracy” metric to ensure consistency across all benchmarks. To ensure more reliable pass@$k$ results, we utilized the estimation method described in [43]. We generated $n$ samples per task, where $n$ is greater than $k$, and evaluated the number of correct samples that passed unit tests. We then calculated the unbiased estimator for pass@$k$. For the self-consistency (Self-Cons.) and verifier results, we randomly selected $k$ out of $n$ samples and conducted separate calculations. The results are reported with an accuracy of ±0.1 at a 95% confidence level. Further details are provided in Section F.2.

### v2

@@ Abstract
In this work, we propose a novel method named Automated Process Labeling via Confidence Variation (AutoCV) to enhance the reasoning capabilities of large language models (LLMs) by automatically annotating the reasoning steps. Our approach begins by training a verification model on the correctness of final answers, enabling it to generate automatic process annotations. This verification model assigns a confidence score to each reasoning step, indicating the probability of arriving at the correct final answer from that point onward. We detect relative changes in the verification’s confidence scores across reasoning steps to automatically annotate the reasoning process. This alleviates the need for numerous manual annotations or the high computational costs associated with model-induced annotation approaches. We experimentally validate that the confidence variations learned by the verification model trained on the final answer correctness can effectively identify errors in the reasoning steps. Subsequently, we demonstrate that the process annotations generated by AutoCV can remarkably enhance the accuracy of the verification model in selecting the correct answer from multiple outputs generated by LLMs. Notably, we achieve substantial improvements across five datasets in mathematics and commonsense reasoning. The source code of AutoCV is available at https://github.com/rookie-joe/AUTOCV.
@@ 3 Method
In the following, we first introduce the problem setting in Section 3.1. Then we discuss the motivation behind why we believe it is necessary to train a verification model in Section 3.2. Finally, we describe how we transition from outcome supervision to process supervision during the training of the verification model in Section 3.3.
@@ 3.2 Motivation
Table 1: Performance of Mixtral-Instruct on the GSM8K test set. All results are reported in %.
@@
Table 2: Comparison of different selection methods across various model sizes (from 7B to 72B) for selecting a response from candidate responses generated by Mixtral-Instruct. All results are reported in %.
@@ 4.1 Experiment on Outcome Supervised Verifier Performance
To evaluate the OSV, we measure its ability to select a sample with the correct final answer from samples generated by various LLMs, denoted as the Response Generator. Specifically, we assess the effectiveness of outcome supervision on two models: Phi2 (OSV-Phi[36] and Mistral-7B (OSV-Mistral[33]. To explore the scalability of this outcome-supervised verifier effect, we choose response generators of varying scales, ranging from 7B to 72B parameters, i.e., Mistral-7B-Instruct (Mistral-7B[33], Mixtral 8 $\times$ 7B (Mixtral-8$\times$7B[32] and Qwen-72B-Chat (Qwen-72B[35]. We employ the metric pass@$k$ [37] to assess the performance of those response generators, defined as the scenario where at least one correct instance is within the model’s first $k$ attempts. We evaluate the OSV’s generalized selection capability across different LLM scales.
Table 3: Performance of OSV models across different configurations. The selection of trained OSV models is based on 5 outputs from response generators.
@@
The performance disparity among the verifiers can be primarily attributed to variations in model sizes and the accuracy of their training data. Specifically, the limited selection capability of the OSV-Phi model stems from discrepancies in the accuracy of its training data. Table Table 4 presents a consolidated view of the model sizes alongside the accuracy metrics of their outcome supervision training data.
@@
In our content, we select OSV-Mistral as the OSV model among other experiment settings due to its superior performance, as demonstrated in Table ??????3.
@@ 5.1 Experimental Setup
Evaluation: We follow evaluation metrics in [42] to ensure consistency across all benchmarks. To ensure more reliable pass@$k$ results, we utilized the estimation method described in [43]. We generated $n$ samples per task, where $n$ is greater than $k$, and evaluated the number of correct samples that passed unit tests. We then calculated the unbiased estimator for pass@$k$. For the self-consistency (Self-Cons.) and verifier’s selection results, we randomly choose $k$ out of $n$ samples and conducted separate calculations. The results are reported with an accuracy of ±0.1 at a 95% confidence level. Further details are provided in Section F.2.